OpenStudy (anonymous):

what is x=12+y and 4x+9y=-17

OpenStudy (anonymous):

Here you go Solve x=12+y and 4x+9y=−17 Steps: I will try to solve your system of equations. x=y+12;4x+9y=−17 Step: Solve x=y+12 for x: Step: Substitute (y+12) for x in 4x+9y=−17: 4x+9y=−17 4(y+12)+9y=−17 13y+48=−17(Simplify both sides of the equation) 13y+48+−48=−17+−48(Add -48 to both sides) 13y=−65 13y 13 = −65 13 (Divide both sides by 13) y=−5 Step: Substitute (−5) for y in x=y+12: x=y+12 x=−5+12 x=7(Simplify both sides of the equation) Answer: x=7 and y=−5

OpenStudy (anonymous):

thanks youre god at this

OpenStudy (anonymous):

I'm not that good.